Fermi National Laboratory
 

Accelerator Update for March 14, 2007 - March 16, 2007

 

Wednesday March 14

The day shift began with Operations monitoring the Tevatron store 5279 (72.35E30) and the Recycler stash (156.1E10), with the Antiproton Source (Pbar) stack at 28.7E30, with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Operations established beam to the Meson MTest experiment at 8:51 AM.

Operations halted beam to Meson at 11:01 AM and MT6 went to open access.

At 11:11 AM, Booster (RF Station repair) and Transport (QS719 shunt work) were prepared for controlled access.

Linac (LRF4 PA work) was prepared for access at 12:12 PM.

Accesses were completed, and beam returns at 4:08 PM.

Operations reported at 4:08 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations terminated store 5279 at 8:05 PM.

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Thursday March 15

The midnight shift began with the Tevatron preparing for a wet squeeze, with the stash (291.4E10), with Pbar stacking (19.1E10), and with NuMI taking beam.

Operations put the Tev into shot setup at 1:35 AM.

Operations established store 5283 at 3:25 AM with an initial luminosity of 197.12E30.

Operations reported at 8:00 AM that the Recycler had stashed antiprotons twice during the midnight shift.

Operations established beam to the Meson MTest experiment at 9:15 AM.

At 3:15 PM a chiller for ECool tripped off. The trip was due to a water leak on the filter unit, experts are replacing the unit.

Operations reported at 4:00 PM that the Recycler had stashed antiprotons three times during the day shift.

MTest turned off at 6:05 PM.

ECool was repaired by 9:31 PM.

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Friday March 16

The midnight shift began with Operations monitoring store 5283 (45.59E30) and the stash (321.7E10), with Pbar stacking (16.8E10), and with NuMI taking beam.

Operations terminated store 5283 at 5:04 AM.

Operations put the Tev into shot setup at 5:19 AM.

Operations established store 5284 at 7:31 AM with an initial luminosity of 196.21E30.

The Plan for the weekend

Stack, stash and store all weekend.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/16/2007    email Fermilab